OpenStudy (anonymous):

Ok Father and Dauther need help with the following x2 + 8x = 180 by completing the square. Thank in advance

OpenStudy (anonymous):

i have my teachers lesson in front of me and the homework answer too; which is -18, 11i don't know how they got that

OpenStudy (amistre64):

to complete a square we simply have to add a useful add the square of half the x coeff to each side of the equation 8/2 = 4; and 4^2 = 16

OpenStudy (anonymous):

the thing to remember for completing the square is (b/2)^2 b over 2 squared you need to add that to each side....awww amistre beat me too it but b is the second term in the trinomial in the form a^2 + b

OpenStudy (amistre64):

x^2 + 8x + 16 = 180 + 16 compact the square (x+4)^2 = 180 + 16 sqrt each side x+4 = +- sqrt(176) subtract the 4 x = -4 +- sqrt(176)

OpenStudy (anonymous):

the answer should be 10 or -18

OpenStudy (amistre64):

180 + 16 = 196 lol

OpenStudy (anonymous):

not 11i

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

you can check the answer by solving the equation on the calculator ;)

OpenStudy (anonymous):

but then how would u solve a similar eqaution tht =0 ?